5. a)consider the following securities and possible state of the economy;
State of the economy Probability Security A Security B Security C Security D
Normal 0.3 0.12 15% 11% 0.130
Boom 0.3 15% 0.110% 0.13 16%
Recession 0.4 17% 12% 11% 20%
Which security will you advise an investor to opt for and why?
b) An investor anticipates Newco’s Security will reach $30 by the end of year. Newco’s beta is 1.3. Assume the return on the market is expected to be 16% and risk free rate is 4%. Calculate the expected return of Newco’s share in one year and determine whether the share is undervalued, overvalued or properly valued with a current value of $25.
c) Differentiate between non-diversifiable risk and liquidity risk.
d) As a trustee of a bond holders, provide two covenants you will insist to be stated in the bond contract. Give reasons for your answer.
1
Expert's answer
2020-11-24T17:36:13-0500

a) Mean(Security A)=0.12*0.3+0.15*0.3+0.17*0.4=0.149

Mean(Security B)=0.15*0.3+0.11*0.3+0.12*0.4=0.126

Mean(Security C)=0.11*0.3+0.13*0.3+0.4*0.11=0.116

Mean(Security D)=0.13*0.3+0.16*0.3+0.20*0.4=0.167

Higher average returns are for security D, so I will advise it to buy.

b) E(R)Newco​=4%+1.3(16%−4%)=20%

And the investor anticipates a 5/25*100%=20% return, so security is properly valued.

c)Non-diversifiable risk can be referred to a risk which is common to a whole class of assets or liabilities. Liquidity risk is a financial risk that for a certain period of time a given financial asset, security or commodity cannot be traded quickly enough in the market without impacting the market price.

d)A bond covenant is a legally binding term of agreement between a bond issuer and a bondholder. I will ask to provide financial information (which allow to monitor company activity) and make sure the assets of the company have adequate insurance (prevent accidental loss).